Suppose that the annual rate of return for a common biotechnology stock is normally distributed with a mean of 4.5% and a standard deviation of 4% . Find the probability that the one-year return of this stock will be positive. Round your answer to at least four decimal places.
Answer by stanbon(75887)

Suppose that the annual rate of return for a common biotechnology stock is normally distributed with a mean of 4.5% and a standard deviation of 4% . Find the probability that the one-year return of this stock will be positive. Round your answer to at least four decimal places.
-----------------
z(0) = (0-4.5)/4 = -1.125
------------------------------
P(x> 0) = P(z > -1.125) = normalcdf(-1.125,100) = 0.8697
===========================================================
Cheers,
Stan H.
===========